OpenStudy (anonymous):

The sum of two numbers is 8 less than twice the first number. Their difference is 4 less than twice the second number. What is the larger number?

OpenStudy (anonymous):

x + y = 2x - 8 (1) x - y = 2y - 4 (2) 2x = 2(x + y) - 12 ---> (x + y) = x + 6 (3) Compare (1) and (3) 2x - 8 = x + 6 --> x = (2) gives 3y = ...... --> y = ...